OpenStudy (anonymous):

Maria is debating between two different mortgages for $155,000. She found a 20-year fixed rate loan at 7.35% and 15-year fixed rate loan at the same rate. How much more interest will she pay for the 20-year loan versus the 15-year loan?

ganeshie8 (ganeshie8):

do you have EMI/monethly payment formula ?

OpenStudy (anonymous):

No, ive looked up three different formula's and none of them with the right answer. Im honestly lost at this point.

ganeshie8 (ganeshie8):

its okay, use below formula : \[\Large\text{monthly payment} = \dfrac{P*i(1+i)^n}{(1+i)^n-1}\] where \(P\) = loan amount = \(155,000\) \(i\) = monthly interest rate = \(\dfrac{0.0735}{12}\) \(n\) = number of months = \(20*12\) or \(15*12\)

ganeshie8 (ganeshie8):

find the monthly payment for each type of loan